# A New Class of Geometric Analog Error Correction Codes for Crossbar Based In-Memory Computing

### 04 Mar 2026 v1

@@ IV-A Dual icosahedral codes
The icosahedron has $12$ vertices, $20$ faces, and $30$ edges. By placing its vertices on the sphere, we obtain $6$ symmetric axes $\{\bm{g}_{\bm{1}},\ldots,\bm{g}_{\bm{6}}\}$, where each axis corresponds to a pair of antipodal vertices. These six axes can be represented by the following matrix:
@@
When searching for the optimal information vector $\bm{u}$} that attains the $m$-height, only its direction matters (i.e., its magnitude does not affect the $m$-height.) Therefore, it is sufficient to restrict the analysis to the faces of the solid, which cover all possible directions. Among the $20$ faces, all are symmetric, and furthermore, each face contains subregions that are themselves symmetric. Thus, analyzing a single representative subregion is equivalent to analyzing the entire space.
@@ IV-B Dual dodecahedral codes
We now consider the geometric code derived from the regular dodecahedron. The dodecahedron has $20$ vertices, $12$ faces, and $30$ edges. Placing its vertices on the sphere yields $10$ symmetric axes $\{{\bm{g}}_{1},\ldots,{\bm{g}}_{10}\}$ (each axis connects a pair of antipodal vertices), represented by
@@
Hence $\beta_{2}$ cannot be the largest among $\{\beta_{2},\beta_{4},\beta_{6},\beta_{7}\}$ since $\beta_{6}\geq\beta_{2}$, and $\beta_{4}$ cannot be the largest since $\beta_{6}\geq\beta_{4}$; therefore the largest must be attained at index $6$ or $7$, i.e., $k=4:\{6,7\}$. Moreover, since both $\beta_{6}$ and $\beta_{7}$ dominate $\beta_{4}$, index $4$ cannot be the second-largest, so the second-largest must lie in $\{2,6,7\}$, i.e., $k=5:\{2,6,7\}$. Next, because $\beta_{6}\geq\beta_{2}$ and $\beta_{6}\geq\beta_{4}$, index $6$ cannot be the third- or fourth-largest within this subset; thus the third-largest must lie in $\{2,4,7\}$, i.e., $k=6:\{2,4,7\}$. Finally, since $\beta_{7}\geq\beta_{4}$, index $7$ cannot be the smallest among the four, so the smallest must lie in $\{2,4\}$, i.e., $k=7:\{2,4\}$.
Within the bottom group $\{3,8,10\}$, inequality (4) gives $\beta_{3}\leq\beta_{10}$, so index $3$ cannot be the largest among $\{3,8,10\}$ and index $10$ cannot be the smallest among $\{3,8,10\}$. Hence
@@
Case I: $m=3,4$ (the denominator is in $\{{\bm{g}}_{2},{\bm{g}}_{6},{\bm{g}}_{7}\}$). By Corollary IV.4, for $m=3,4$, the (m+1)(m\!+\!1)-th largest projection must be attained at index $2$, 66, or $7$.
@@
For $f_{2}$, both $\partial_{u}f_{2}$ and $\partial_{v}f_{2}$ are strictly positive. Evaluating $f_{2}$ along edge $AB$ shows that
@@
so $f_{2}$ is strictly increasing along $AB$. Hence the maximum on $AB$ is attained at $u=1$, i.e., at the vertex $A=(1,0)$.
For $f_{6}$, both $\partial_{u}f_{6}$ and $\partial_{v}f_{6}$ are strictly positive. Evaluating $f_{6}$ along edge $AB$ shows that
@@
so $f_{6}(u,1-u)$ is strictly increasing on $u$. Therefore the maximum on edge $AB$ is attained at $u=1$ (i.e., $(u,v)=(1,0)$, the vertex $A$).
@@
Case II: $m=5,6$ (the denominator is in $\{{\bm{g}}_{2},{\bm{g}}_{4},{\bm{g}}_{7}\}$). For $m=5$ or $m=6$, the (m+1)(m\!+\!1)-th largest magnitude must be attained at index $2$, 44, or $7$. On $T^{\prime}$, ${\bm{x}}\cdot{\bm{g}}_{2}\geq 0$ and ${\bm{x}}\cdot{\bm{g}}_{7}\geq 0$, while ${\bm{x}}\cdot{\bm{g}}_{4}<0$, so $|{\bm{x}}\cdot{\bm{g}}_{2}|={\bm{x}}\cdot{\bm{g}}_{2}$, $|{\bm{x}}\cdot{\bm{g}}_{7}|={\bm{x}}\cdot{\bm{g}}_{7}$, and $|{\bm{x}}\cdot{\bm{g}}_{4}|=-({\bm{x}}\cdot{\bm{g}}_{4})$. The switching interfaces are
@@
Therefore, the maximum is attained at the endpoint with the largest $u$, i.e., at the vertex $A$.
For $f_{4}$, we have $\partial_{v}f_{4}<0$. Hence any maximizer must lie on $CA\cap T_{2}$ or on $L_{2,7}$. From the previous calculation, along $CA$ we also have $\partial_{u}f_{4}<0$. Therefore, if the maximizer lies on $CA\cap T_{2}$, it must occur at the endpoint where $CA\cap T_{2}$ meets $L_{2,7}$. Consequently, it suffices to restrict attention to $L_{2,7}$. Evaluating $f_{4}$ along edge $L_{2,7}$ shows that
@@
The maximuzer for $f_{2}$ in $T_{3}$ is on edge $AB$, maximizing $f_{2}$ along $AB$ is already done in previous case.
@@
Therefore, the maximum of $f_{10}$ over $T_{1}$ is attained at the vertex with minimal $u$ and maximal $v$, namely at the intersection of $L_{8,10}^{-}$ and the edge $BC$, which is

### 10 Apr 2026 v3

@@ IV-A Dual icosahedral codes
The icosahedron has $12$ vertices, 2020 faces, and $30$ edges. By placing its vertices on the sphere, we obtain 66 symmetric axes $\{\bm{g}_{\bm{1}},\ldots,\bm{g}_{\bm{6}}\}$, where each axis corresponds to a pair of antipodal vertices. These six axes can be represented by the following matrix:
@@
When searching for the optimal information vector $\bm{u}$} that attains the $m$-height, only its direction matters (i.e., its magnitude does not affect the $m$-height.) Therefore, it is sufficient to restrict the analysis to the faces of the solid, which cover all possible directions. Among the 2020 faces, all are symmetric, and furthermore, each face contains subregions that are themselves symmetric. Thus, analyzing a single representative subregion is equivalent to analyzing the entire space.
@@ IV-B Dual dodecahedral codes
We now consider the geometric code derived from the regular dodecahedron. The dodecahedron has 2020 vertices, $12$ faces, and $30$ edges. Placing its vertices on the sphere yields $10$ symmetric axes $\{{\bm{g}}_{1},\ldots,{\bm{g}}_{10}\}$ (each axis connects a pair of antipodal vertices), represented by
@@
Hence $\beta_{2}$ cannot be the largest among $\{\beta_{2},\beta_{4},\beta_{6},\beta_{7}\}$ since $\beta_{6}\geq\beta_{2}$, and $\beta_{4}$ cannot be the largest since $\beta_{6}\geq\beta_{4}$; therefore the largest must be attained at index 66 or 77, i.e., $k=4:\{6,7\}$. Moreover, since both $\beta_{6}$ and $\beta_{7}$ dominate $\beta_{4}$, index 44 cannot be the second-largest, so the second-largest must lie in $\{2,6,7\}$, i.e., $k=5:\{2,6,7\}$. Next, because $\beta_{6}\geq\beta_{2}$ and $\beta_{6}\geq\beta_{4}$, index 66 cannot be the third- or fourth-largest within this subset; thus the third-largest must lie in $\{2,4,7\}$, i.e., $k=6:\{2,4,7\}$. Finally, since $\beta_{7}\geq\beta_{4}$, index 77 cannot be the smallest among the four, so the smallest must lie in $\{2,4\}$, i.e., $k=7:\{2,4\}$.
Within the bottom group $\{3,8,10\}$, inequality (4) gives $\beta_{3}\leq\beta_{10}$, so index 33 cannot be the largest among $\{3,8,10\}$ and index $10$ cannot be the smallest among $\{3,8,10\}$. Hence
@@
Case I: $m=3,4$ (the denominator is in $\{{\bm{g}}_{2},{\bm{g}}_{6},{\bm{g}}_{7}\}$). By Corollary IV.4, for $m=3,4$, the (m+1)(m\!+\!1)-th largest projection must be attained at index 22, 66, or 77.
@@
For $f_{2}$, both $\partial_{u}f_{2}$ and $\partial_{v}f_{2}$ are strictly positive. Evaluating $f_{2}$ along edge ABAB shows that
@@
so $f_{2}$ is strictly increasing along ABAB. Hence the maximum on ABAB is attained at $u=1$, i.e., at the vertex $A=(1,0)$.
For $f_{6}$, both $\partial_{u}f_{6}$ and $\partial_{v}f_{6}$ are strictly positive. Evaluating $f_{6}$ along edge ABAB shows that
@@
so $f_{6}(u,1-u)$ is strictly increasing on $u$. Therefore the maximum on edge ABAB is attained at $u=1$ (i.e., $(u,v)=(1,0)$, the vertex AA).
@@
Case II: $m=5,6$ (the denominator is in $\{{\bm{g}}_{2},{\bm{g}}_{4},{\bm{g}}_{7}\}$). For $m=5$ or $m=6$, the (m+1)(m\!+\!1)-th largest magnitude must be attained at index 22, 44, or 77. On $T^{\prime}$, ${\bm{x}}\cdot{\bm{g}}_{2}\geq 0$ and ${\bm{x}}\cdot{\bm{g}}_{7}\geq 0$, while ${\bm{x}}\cdot{\bm{g}}_{4}<0$, so $|{\bm{x}}\cdot{\bm{g}}_{2}|={\bm{x}}\cdot{\bm{g}}_{2}$, $|{\bm{x}}\cdot{\bm{g}}_{7}|={\bm{x}}\cdot{\bm{g}}_{7}$, and $|{\bm{x}}\cdot{\bm{g}}_{4}|=-({\bm{x}}\cdot{\bm{g}}_{4})$. The switching interfaces are
@@
Therefore, the maximum is attained at the endpoint with the largest $u$, i.e., at the vertex AA.
For $f_{4}$, we have $\partial_{v}f_{4}<0$. Hence any maximizer must lie on $CA\cap T_{2}$ or on $L_{2,7}$. From the previous calculation, along CACA we also have $\partial_{u}f_{4}<0$. Therefore, if the maximizer lies on $CA\cap T_{2}$, it must occur at the endpoint where $CA\cap T_{2}$ meets $L_{2,7}$. Consequently, it suffices to restrict attention to $L_{2,7}$. Evaluating $f_{4}$ along edge $L_{2,7}$ shows that
@@
The maximuzer for $f_{2}$ in $T_{3}$ is on edge ABAB, maximizing $f_{2}$ along ABAB is already done in previous case.
@@
Therefore, the maximum of $f_{10}$ over $T_{1}$ is attained at the vertex with minimal $u$ and maximal $v$, namely at the intersection of $L_{8,10}^{-}$ and the edge BCBC, which is